OpenStudy (anonymous):

f(x) = Square root of quantity x plus seven.; g(x) = 8x - 11 Find f(g(x)).

OpenStudy (ivettef365):

so you have f(x) = \(\sqrt{x+7}\) and g(x)= 8x-11

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes.

OpenStudy (ivettef365):

the only thing you need to do is substitute the x from f(x) for the value of g(x) which is 8x -11

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

I still don't get it...

OpenStudy (ivettef365):

you have f(x) = \(\sqrt{(8x-11)+7}\) you see how I substitute the x

OpenStudy (ivettef365):

now that is \(\sqrt{8x-4}\)

terenzreignz (terenzreignz):

Here's an example: Let's have \[\Large p(\color{red}x) = 8\color{red}x-4\] and \[\Large q(x) = \color{blue}{3x+1}\] So to get \[\Large p(q(x))\]You just replace the \(\large \color{red}x\) in \(\large p(\color{red}x)\) with \(\large q(x) \)itself like so: \[\Large p(q(x))= 8(\color{blue}{3x+1})-4\] and then simplify: \[\Large p(q(x)) = 24x +4\]

OpenStudy (anonymous):

@ivettef365 would the final solution be f(g(x)) = . @terenzreignz Thank You for that example, made things a bit simple.